OpenStudy (love_to_love_you):

Write a recursive definition. Show all work. 80, 77, 74, 71, 68...

OpenStudy (love_to_love_you):

@AlexandervonHumboldt2

OpenStudy (love_to_love_you):

@ganeshie8

OpenStudy (love_to_love_you):

@xapproachesinfinity

OpenStudy (anonymous):

x-3

OpenStudy (love_to_love_you):

is that it?

OpenStudy (love_to_love_you):

how did you get that

OpenStudy (love_to_love_you):

@AJ01 please hurry I have a deadline

OpenStudy (anonymous):

n1=n-3 n1=(80)-3=77 n=(77)-3=74